OpenStudy (anonymous):

2sin²3Ө+3sin3Ө+1=0 Find all degree solutions.

OpenStudy (anonymous):

let \[\sin3 \theta = x\] what does that equation become?

OpenStudy (anonymous):

2x^2+3x+1=0

OpenStudy (anonymous):

can you solve that for x?

OpenStudy (anonymous):

can i use quadratic formula for that?

OpenStudy (anonymous):

if you want, or you can factorise it

OpenStudy (anonymous):

x=-1/2 and x=-1

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

now change those solutions back to trig using what we did we had earlier \[x = \sin 3 \theta\]

OpenStudy (anonymous):

\[\sin 3 \theta = \frac{-1}{2} \text{ , } -1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what else ?

OpenStudy (kinggeorge):

Now you would need to solve the \(\theta\) in each equation using the unit circle.

OpenStudy (anonymous):

is there any other way other than the unit circle ?

OpenStudy (anonymous):

is 70 degrees and 90 degrees are answers ?

OpenStudy (kinggeorge):

The unit circle is the best way, but there are other methods.

OpenStudy (anonymous):

2sin²3Ө+3sin3Ө+1=0 Let sin3Ө=x so 2x^2+3x+1=0 (x+1)(2x+1)=0 so x=-1 or x=-1/2 so sin3Ө=-1 or sin3Ө=-1/2 3Ө=-90 +k360 = -30+k120 or 3Ө=270+k360=90 + k120 for sin3Ө=-1 or 3Ө=-30+k360=-10+k120 or 3Ө=210+k360=70+k120 for sin3Ө=-1/2 where k is integer

OpenStudy (anonymous):

yeah i got 70 and 90 just like Keroro :|

OpenStudy (anonymous):

@Keroro u think it's right ?

OpenStudy (anonymous):

yip :) if this is high school level, its definitely correct

OpenStudy (anonymous):

so there are only 2 answers ?

OpenStudy (anonymous):

i thought on problems like this, always like 3 or 4 answers :|

OpenStudy (anonymous):

Assuming that the question is to find the general solution, yes, it is only 2 answers but you must include the periods i.e. k360 etc

OpenStudy (anonymous):

which means that there can be plenty of solutions but you know u are on the right track because you have "squared" in your question which means that at most you must have 2 answers